# Solving Distributed Flexible Job Shop Scheduling Problems in the Wool Textile Industry with Quantum Annealing

### v1

@@ 4 Modell
In this Section we describe the QUBO model to solve the extended DFJSP as required by the wool textile industry. To this end, the extended problem is first formally described in Subsection 4.1. Based on this, we create the QUBO model in Subsection 4.2, reduce the number of its variables in Subsection 4.3 and deal with the determination of the Lagrange parameters in Subsection 4.4.
@@ 4.1 Problem formulation
The DFJSP refers to job scheduling in distributed manufacturing environments. There could be different objectives when solving the DFJSP (e.g., minimizing job completion time, processing costs or energy consumption). This paper focuses on finding the optimal schedule that minimizes the time required to complete the jobs, i.e., the makespan. The formulation used is the one introduced by Venturelli et al. [24] for the JSSP and the extension from Schworm et al. [21] for the FJSP. In this research, the model is extended to include shipping time for distributed operations. It is assumed that there are A jobs J = {j, …, j} to be processed on B machines M = {m, …, m} in C factories F = {f, …, f}. Each job consists of a set of L operations = {o, …, o}. Any operation can be processed on at least one machine out of the given set M. For any job, the order of operations is given by = {r, …, r}, in which r is the last operation of job . Another notations used in this paper are listed in Tbl. 1.
@@
| o | Next operation of any operation . |
| m | Machine of next operation o. |
@@
| t | Next time step of any time step . |
@@
Reminder: The production orders are referred to as jobs and the production steps as operations in the context of the DFJSP. Since machines are associated with factories, the distance between machines is derived from the distance between factories. Distance between factories is directly related to the shipping time needed to send the lots in production to another location. In this paper, distance and shipping time is used interchangeably, and refer to the value that is added to the cost function when production lots are shipped between locations. Fig. 2 shows a schema representing the DFJSP.
Different jobs may have a different number of operations, each with its own processing time on a specific machine. This paper focuses on a special case of DFJSP where operations of the same job can be performed on machines in geographically distributed factories, which means that the distance between machines should be considered when generating the job schedule.
@@
Precedence of operations - an operation can only start, when the previous operation has finished.
@@
Operation start once - operation processing on machines cannot be interrupted (the reason behind this condition is that each machine must be technically adjusted for every new production lot, which is time-consuming).
@@
Other underlying assumptions of the DFJSP are that operations of a job must be performed in the correct order to complete that job. The operations of different jobs are not subject to any precedence constraints and can be processed in parallel. Each factory is equipped with machines of different types, thus leading to different potential processing times of an operation. A job can be scheduled on a machine more than once. All machines are available for use and all jobs can be processed starting with time zero [28].
@@ 4.2 QUBO model
Samplers are processes used to solve problems with D-Wave. A BQM sampler samples low energy states in the model defined by the QUBO equation and returns a list of samples to increase energy. The QUBO model is composed of binary variables x, each represent one operation per unit of time on a machine, up to the makespan:
@@
The objective function H(x) is defined as a cost function (Eq. 5), which minimizes the time taken to complete all operations (the makespan). Cost functions are mathematical expressions of optimization problems. For quantum computing, cost functions are quadratic models that have the lowest values (energy) for good solutions to the problems they represent. Each point on a cost function represents a different solution configuration - in this case, each configuration is a particular assignment of the starting times of operations. The goal of the optimization is to minimize the cost of the solution, i.e. the amount of time needed to complete all operations. The mathematical representation of the penalty terms and of the makespan objective are added and build the cost function of the DFJSP with the format:
@@
f(x), g(x) and h(x) represent penalty functions for precedence, operation once and no-overlap constraints, respectively,
@@
Every time a solution violates one or more constraints, a penalty is applied to that solution (e.g., two operations on a single machine that are scheduled to run at the same time violate the no-overlap constraint, so a penalty will be assigned). Positive terms penalize certain solution configurations, while negative ones support them. By adding penalties to terms that break the constraints, the cost of those configurations increases. This reduces the likelihood that a suboptimal solution will be generated. The importance attributed to each term can be adjusted using the weights , and . The process of adjusting these weights is explained in Section 4.4. The weights determine the importance of penalty functions in relation to other terms.
Precedence constraint. The precedence constraint states that a job operation can only start after the previous operation is finished and the lot arrived at the production site of the next operation if shipping is necessary.
To formulate the penalty, cases where consecutive operations o and o in a job take place out of order are counted. If the starting time of o denoted by t is less than the sum of the starting time of o denoted by t, its processing time p and the shipping time between machines d, then this count as penalty. The penalty condition can be formulated as follows:
@@
represent the expression that assigns a penalty every time the precedence constraint is violated.
@@
represent the expression that assigns a penalty every time the precedence constraint is violated,
@@
The set A is a condition that forbids operation o to start at time t if the processing of the previous operation o has not finished yet (meaning that ). The set B forbids the operations to start at the same time. Further details on the penalty function are:
@@
Makespan Function. To formulate the optimization objective of completing each job in the shortest time possible, the following polynomial function is defined, that penalizes the late completion time of operations:
@@ 4.3 Variable pruning
In the previous section, the binary variable x was defined to represent whether an operation o of a job i is scheduled to start on a machine m at time t or not. To reduce the number of variables in the model, a method called variable pruning is used, which removes variables that are not possible in valid solutions. First, an upper bound t of the highest possible makespan is calculated. Without this step, the solver will have to consider all solutions between zero to infinity, which implies a very large solution space that makes it difficult to embed and solve the QUBO [24]. Instead, the search space is limited by defining the upper bound of a schedule as the sum of the maximum processing time of each operation in all jobs and maximum shipping time between these operations. Thus, a significant number of variables x can be pruned by applying restrictions on time t.
Next, the duration and order of operations within a job are examined. Variables that allow an operation to execute at a given time that results in an invalid schedule are removed. Specifically, variables are pruned to allow operations to start earlier than the sum of the processing time of its predecessors, or if they allow operations to start later than the sum of its own processing time and that of the successor operations subtracted from t [29].
@@
P is the minimum predecessor time of the operation o, calculated as the sum of the minimum processing time of the preceding operations and shipping distance between them. A starting time earlier than the minimum predecessor time is invalid, therefore, these binary variables can be eliminated from the model [21].
@@
These variables can be also eliminated from the QUBO, since they represent invalid starting times [24]. Thus, the starting time of operations o should be in the range , all the other variables that are not in this range are removed from the model to decrease the size of the computational problem and computation time [29].
@@ 4.4 Lagrange parameters
In the previous sections the constraints have been formulated as binary polynomials in a Hamilton formulation. The binary polynomials f(x), g(x), h(x) are multiplied with non-negative scalar weights also known as Lagrange parameters (Eq. 5), which determine the impact of the respective polynomial [21]. Weights are assigned to each constraint depending on the relative strengths of the objective function and the constraints. The weights have to be set high enough, so that valid solution with no broken constraints will be returned. On the other hand, values that are too high will result in solutions that are valid, but far from optimal777D-Wave Problem Formulation Guide Whitepaper: https://www.dwavesys.com/media/bu0lh5ee/problem-formulation-guide-2022-01-10.pdf.
@@
For example, for job j in Tbl. 2, operation o is starting on machine m at time t and has a processing time p (this technique is applied in the same way for larger problem instances with more jobs and operations).
@@
Thus, the sum of processing times of previous operations of o is equal to 1. Further, o of the same job is starting at different times t on machine m. Shipping time between machines d equals to zero. Both binary variables x and x are equal to 1. This can be expressed by a QUBO. For , the value and thus, the energy of cost function H(x) equals to 3:
This is the sum of the processing times of the two operations o and o that represent the makespan objective function (second term in equation H(x)). Since the start-once constraint is not violated, the first term in equation H(x) is equal to zero. This is also the minimum energy level of a valid solution for this specific problem instance. Any violation of the constraints will add penalties to the cost function that will increase the energy of the system.
A specific case should, however, be considered. Let us assume that operation o is not scheduled to start at all. This will result in a cost function H(x) that equals to 2:
To compensate for this effect, a Lagrange parameter is applied that increases the weight of the penalty term. In DFJSP valid solutions with higher energy are preferred to invalid solutions with low energy but violated constraints. The value of the weight parameter depends on the problem instance. A value that is too low penalizes invalid configurations of variables insufficiently, resulting energy values that might be lower than the minimum energy of the best valid solution. Values that are too high would return valid configurations – possible with higher energy than the optimum. The value of the Lagrange parameter is set based on the maximum makespan of the jobs ( t in Section 4.3). The maximum energy value, which corresponds to the maximum makespan t, is equal to 10 (this value is calculated for each problem instance based on the input job dictionary). The penalty is set to , meaning the maximum energy possible for a valid configuration for this problem instance, which results in the following function H(x):
The energy of the function is equal to , which is above the maximum energy possible for valid solutions of . This ensures that no invalid solution will have lower energy than any of the valid solutions. The technique demonstrated in this simplified example can be applied in the same way to problems that have more jobs and operations. The resulting QUBO will have more variables.
The same method is used to calculate the weights for other constraints. Generally, the Lagrange parameters are set individually for each constraint, depending on the magnitude of the objective function to be minimized and its relative strength to the constraints. In this case, to make sure that the cost function H(x) has the lowest energy for valid solutions, the weights of penalty functions f(x), g(x) and f(x) are all set equal to t, which is the maximum energy value of all valid solutions.
@@ 5 Experiments
In this Section, we implement the model and conduct experiments on a quantum annealer of D-Wave (more precisely, the AdvantageSystem4.1). As it is common practice, the results of the experiment on a quantum annealer are compared with results obtained from SA. Firstly, the problem size (see 5.1) and solver parameters (see 5.2) are defined. The QUBO model is constructed with PyQUBO Python library, which is a high-level interface that helps formulating combinatorial optimization problems in QUBO form, and supports solving them using various solvers, including D-Wave quantum annealer [27]. We then present the results of the experiments in Subchapter 5.3 and discuss them in Subchapter 5.4.
@@ 5.1 Problem size and dataset
To analyze the performance of the D-Wave quantum annealer for the DFJSP, 6 problem instances are defined in such a way that the size of the smallest problem instance is 50 and the size of the largest one is 250. The instances in between are defined in even intervals of 50 (e.g. 100, 150, etc888Problem size is rounded for some instances (e.g. instance size 99 is rounded to 100).). The size of the problem instances refers to the number of binary variables x necessary to represent the problem in QUBO form. The maximum size of the problem instances is dictated by the physical properties of the D-Wave QPU. The D-Wave AdvantageSystem4.1 has 5627 qubits. However, not all qubits are interconnected. When embedding the model variables onto the physical hardware graph of the QPU, in some cases, to represent a single logical variable set of physical qubits called chains are necessary, with couplings between physical qubits capable to realize the correct interaction between logical variables [26]. This leads to embedding patterns that need more physical qubits than there are variables in the original problem.
The purpose of the experiment is to test the capacity and performance of the D-Wave quantum annealer for the DFJSP up to the largest problem that can be embedded into the QPU. Tbl. 3 summarizes the instances solved by QA, the number of qubits needed to embed the model onto the physical QPU and the length of the longest chain in the embedding.
@@
Smaller problem instances require a lower number of qubits for embedding. It is also true that the length of the chains increases with problem size, which means that with increasing number of variables, it is more difficult to find embedding schemas. Therefore, longer chains are required to connect the variables through couplers. In this model, the largest problem instance that could be embedded was of size 250, that required more than 4000 qubits for embedding. For 300 variables, no embedding could be found.
The problem instances are formulated for 2 jobs, j and j, each job containing a minimum of 2 operations. Each operation can be processed on one machine or a set of two machines, with different processing times on each machine. To include shipping distance in the model, the sets of machines available for each operation include machines from different factories. The consideration of the shipping distance is essential for the processing of Contribution 1. The number of operations in each job, the available machines and the processing time on each machine determine the number of binary variables x necessary to represent the problem in QUBO form.
@@ 5.2 Solver parameters
There are 20 solver parameters [6] that can be adjusted when submitting problems to solvers. In this experiment, the default settings are kept for most of them, including the default annealing time of 20s. The parameters that are adjusted in the experiment are the number of reads and the chain strength.
Number of reads. In theory, very long annealing times should keep the system in a low energy state, in practice noise due to hardware imperfections will push the state to a higher energy [4]. Therefore, short annealing times are preferred in order to take many samples from the energy distribution of the problem. The appropriate number of samples depends on the problem itself and on its energy distribution and is usually between 10 and 10000 samples [4]. In this experiment, the number of reads (’num_reads’) is set to 1000. Generally, increasing the number of reads can increase the probability to get better solutions, but improvements diminish due to some problem-dependent numbers [6]. For this model, increasing the number of reads much above 1000 does not seem to be of any advantage. Sampling the problem instance with 200 variables 3000 times produced a solution with an energy of 34 and 2 violated constraints, which is worse than the energy of the best solution of 24 and 1 violated constraint obtained with 1000 reads. This is consistent with the results obtained by Carugno et al. [4], who found that increasing the number of reads from 1000 to 10000 did not have any beneficial effects.
Chain strength. For a solution to be consistent, the chain of qubits has to act as a single qubit, meaning that all the qubits in the chain must have the same values after annealing [4]. Broken chains show different values of the qubits at the end of the annealing process. Chains must be set strong enough to ensure that the qubits of the chain are showing the same value compared to other qubits. If the chains strength is too small, then the qubits will show different values. If the chain strength parameter is too high, this will impede the qubits in the chain to change their state, the chain begins to behave like a separate entity that does not interact with other terms of the function.
We follow the D-Wave recommendation to determine the chain strength for each problem instance and start with two utility functions to estimate the chain strength [7]:
@@
The first function is returning a chain strength that is scaled to the problem bias range. The second function calculates a chain strength that attempts to compensate for the torque that would break the chain. Both methods are to be seen as a starting point to estimate the optimal chain strength, and not as rule, since this parameter is problem-dependent. In the next step, the chain strength parameter found in step 1 should be doubled or tripled, to check if this change has an impact on the percentage of broken chains.
@@
To find the best chain strength for the model, the model is run on the QPU for each problem instance, and the resulting energy and the percentage of broken chains in the best solution are analyzed. The intention is to find the lowest energy possible, to at the same time having a low percentage of broken chains to ensure a consistent solution. Fig. 3 shows the results of the measurements for the problem instance of size 150.
@@
The utility function scaled returns a chain strength of 32 with an energy of 69, the function uniform_torque returns a value of 309 with an energy of 966. To understand the relation between these parameters, the range between these two chain strengths is divided into 11 intervals and the energy and the percentage of broken chains in the best solutions for each of these points is measured. Additionally, the smallest chain strength value is divided by 2, to check whether a smaller parameter could improve the results. The results in the graph show that the lowest energy achieved corresponds to the next point in the interval from the point obtained with the scaled function. However, the percentage of broken chains corresponding to this value is 41%, which is high compared to values in the graph for higher chain strength values. In order to decide whether to accept a chain strength parameter of 39 or not, an additional analysis of the consistency of the solution is necessary. Fig. 4 shows that, despite a higher number of broken chains, the embedding with a chain strength of 39 delivers a consistent solution without broken constraints. With the increasing chain strength parameter the number of broken constraints increases too.
@@
This procedure is repeated in the experiment for all problem instances to calculate the appropriate parameter for the chain strength. For all instances, the parameter that delivered the best solution with the lowest energy is one interval above the parameter estimated by the scaled function. Since the intention is to test to which extent the QPU is able to handle large problems, this method is applied to problem instances of size 250 and 300. For problems of size 50 to 150, the solver is able to find consistent solutions with no broken constraints. For problem instances of size 200 and 250, the solutions with the lowest energy found had 1 and 2 broken constraints, respectively. For the instance of size 300 no embedding could be found.
@@
SA parameters. To solve problems with SA, the D-Wave Sampler class SimulatedAnnealingSampler of the D-Wave Ocean SDK is used. The sampler utilizes a classical heuristic SA algorithm that runs locally on the CPU999Hardware configuration: 2x Intel Xeon Gold 6326 CPU 16Core, 2.9GHz 24M 185W Socket 4189, 4 x Samsung 64GB DDR4 ECC RDIMM 3200Mhz total 256GB RAM, 4 x HDD WD Gold Datacenter 1TB SATA, 1x SSD Samsung PM883 480GB SATA, 2 x 1 / 10 GB Ethernet 10 BaseT RJ-45 onboard.. For SA, the Central Processing Unit (CPU) time is measured, which is the amount of time that the CPU has spent solving the problem. The CPU time is compared to QPU access time.

### v3

@@ 4 Modell
In this Section, we describe the QUBO model for solving the extended DFJSP as required by the wool textile industry. To this end, the extended problem is first formally described in Subsection 4.1. Based on this, we create the QUBO model in Subsection 4.2, reduce the number of its variables in Subsection 4.3, and deal with the determination of the Lagrange parameters in Subsection 4.4.
@@ 4.1 Problem formulation
The DFJSP refers to job scheduling in distributed manufacturing environments. There could be different objectives when solving the DFJSP (e.g., minimizing job completion time, processing costs, or energy consumption). This paper focuses on finding the optimal schedule that minimizes the time required to complete the jobs, i.e., the makespan. The formulation used is the one introduced by Venturelli et al. [33] for the JSSP and the extension from Schworm et al. [25] for the FJSP. In this research, the model is extended to include shipping time for distributed operations. It is assumed that there are A jobs J = {j, …, j} to be processed on B machines M = {m, …, m} in C factories F = {f, …, f}. Each job consists of a set of L operations = {o, …, o}. Any operation can be processed on at least one machine out of the given set M. For any job, the order of operations is given by = {r, …, r}, in which r is the last operation of job . Another notations used in this paper are listed in Tbl. 1.
@@
| o′ | Next operation of any operation . |
| m′ | Machine of next operation o. |
@@
| t′ | Next time step of any time step . |
@@
Reminder: The production orders are referred to as jobs and the production steps as operations in the context of the DFJSP. Since machines are associated with factories, the distance between machines is derived from the distance between factories. Distance between factories is directly related to the shipping time needed to send the lots in production to another location. In this paper, ’distance’ and ’shipping time’ are used interchangeably and refer to the value added to the cost function when production lots are shipped between locations. Fig. 2 shows a schema representing the DFJSP.
Different jobs may have varying numbers of operations, each with its own processing time on a specific machine. This paper focuses on a special case of DFJSP where operations of the same job can be performed on machines in geographically distributed factories, which means that the distance between machines should be considered when generating the job schedule.
@@
Precedence of operations - an operation can only start when the previous operation has finished.
@@
Operation starts once - operation processing on machines cannot be interrupted (the reason ,behind this condition is that each machine must be technically adjusted for every new production lot, which is time-consuming).
@@
Other underlying assumptions of the DFJSP are that operations of a job must be performed in the correct order to complete that job. The operations of different jobs are not subject to any precedence constraints and can be processed in parallel. Each factory is equipped with machines of various types, resulting in different potential processing times for an operation. A job can be scheduled on a machine multiple times. All machines are available for use and all jobs can be processed starting with time zero [39].
@@ 4.2 QUBO model
Samplers are processes used to solve problems with D-Wave. A BQM sampler samples low-energy states in the model defined by the QUBO equation and returns a list of samples to increase energy. The QUBO model is composed of binary variables x, each representing one operation per unit of time on a machine, up to the makespan:
@@
The objective function H(x) is defined as a cost function (Eq. 5), which minimizes the time taken to complete all operations (the makespan). Cost functions are mathematical expressions of optimization problems. For quantum computing, cost functions are quadratic models that have the lowest values (energy) for good solutions to the problems they represent. Each point on a cost function represents a different solution configuration - in this case, each configuration is a particular assignment of the starting times of operations. The goal of the optimization is to minimize the cost of the solution, i.e., the amount of time needed to complete all operations. The mathematical representation of the penalty terms and the makespan objective are added and build the cost function of the DFJSP with the format:
@@
f(x), g(x), and h(x) represent penalty functions for precedence, operation once, and no-overlap constraints, respectively,
@@
Every time a solution violates one or more constraints, a penalty is applied to that solution (e.g., two operations on a single machine that are scheduled to run at the same time violate the no-overlap constraint, which means that a penalty will be assigned). Positive terms penalize certain solution configurations, while negative ones support them. By adding penalties to terms that break the constraints, the cost of those configurations increases. This reduces the likelihood of generating a suboptimal solution. The importance attributed to each term can be adjusted using the weights , , and . The process of adjusting these weights is explained in Section 4.4. The weights determine the importance of penalty functions in relation to other terms.
Precedence constraint. The precedence constraint states that a job operation can only start after the previous operation is finished and the lot arrived at the production site of the next operation, if shipping is necessary.
To formulate the penalty, cases where consecutive operations o and o′ in a job take place out of order are counted. If the starting time of o′ denoted by t is less than the sum of the starting time of o denoted by t, its processing time p and the shipping time between machines d, then this counts as penalty. The penalty condition can be formulated as follows:
@@
represents the expression that assigns a penalty every time the precedence constraint is violated.
@@
represents the expression that assigns a penalty every time the precedence constraint is violated,
@@
The set Am is a condition that forbids operation o′ to start at time t′ if the processing of the previous operation o has not finished yet (meaning that ). The set Bm forbids the operations to start at the same time. Further details on the penalty function are:
@@
Makespan Function. To formulate the optimization objective of completing each job in the shortest time possible, the following polynomial function is defined, which penalizes the late completion time of operations:
@@ 4.3 Variable pruning
In the previous Section, the binary variable x was defined to represent whether an operation o of a job i is scheduled to start on a machine m at time t or not. To reduce the number of variables in the model, a method called variable pruning is used, which removes variables that are not possible in valid solutions. First, an upper bound t of the highest possible makespan is calculated. Without this step, the solver will have to consider all solutions between zero and infinity, which implies a very large solution space that makes it difficult to embed and solve the QUBO [33]. Instead, the search space is limited by defining the upper bound of a schedule as the sum of the maximum processing time of each operation in all jobs and the maximum shipping time between these operations. Thus, a significant number of variables x can be pruned by applying restrictions on time t.
Next, the duration and order of operations within a job are examined. Variables that prevent an operation from executing at a given time, resulting in an invalid schedule, are removed. Specifically, variables are pruned to allow operations to start earlier than the sum of the processing time of its predecessors, or if they allow operations to begin later than the sum of its own processing time and that of the successor operations subtracted from t [40].
@@
P is the minimum predecessor time of the operation o, calculated as the sum of the minimum processing time of the preceding operations and shipping distance between them. A starting time earlier than the minimum predecessor time is invalid; therefore, these binary variables can be eliminated from the model [25].
@@
These variables can also be eliminated from the QUBO, since they represent invalid starting times [33]. Thus, the starting time of operations o should be in the range , all the other variables that are not in this range are removed from the model to decrease the size of the computational problem and computation time [40].
@@ 4.4 Lagrange parameters
In the previous sections, the constraints have been formulated as binary polynomials in a Hamilton formulation. The binary polynomials f(x), g(x), h(x) are multiplied with non-negative scalar weights also known as Lagrange parameters (Eq. 5), which determine the impact of the respective polynomial [25]. Weights are assigned to each constraint depending on the relative strengths of the objective function and the constraints. The weights must be set high enough so that a valid solution with no broken constraints is returned. On the other hand, values that are too high will result in solutions that are valid, but far from optimal888Problem Formulation Guide [White Paper]: https://www.dwavesys.com/media/bu0lh5ee/problem-formulation-guide-2022-01-10.pdf.
@@
For example, for job j1 in Tbl. 2, operation o1 is starting on machine m at time t and has a processing time p (this technique is applied in the same way for larger problem instances with more jobs and operations).
Thus, the sum of processing times of previous operations of o2 is equal to 1. Further, o2 of the same job is starting at different times t on machine m. Shipping time between machines d equals to zero. Both binary variables x and x are equal to 1. This can be expressed by a QUBO. For , the value and thus, the energy of cost function H(x) equals to 3:
This is the sum of the processing times of the two operations o1 and o2 that represent the makespan objective function (second term in equation H(x)). Since the start-once constraint is not violated, the first term in equation H(x) is equal to zero. This is also the minimum energy level of a valid solution for this specific problem instance. Any violation of the constraints will add penalties to the cost function that will increase the energy of the system.
A specific case, however, should be considered. Let us assume that operation o2 is not scheduled to start at all. This will result in a cost function H(x) that equals to 2:
@@
To compensate for this effect, a Lagrange parameter is applied that increases the weight of the penalty term. In DFJSP, valid solutions with higher energy are preferred to invalid solutions with low energy but violated constraints. The value of the weight parameter depends on the problem instance. A value that is too low penalizes invalid configurations of variables insufficiently, resulting in energy values that might be lower than the minimum energy of the best valid solution. Values that are too high would return valid configurations – possible with higher energy than the optimum. The value of the Lagrange parameter is set based on the maximum makespan of the jobs ( t in Section 4.3). The maximum energy value, which corresponds to the maximum makespan t, is equal to 10 (this value is calculated for each problem instance based on the input job dictionary). The penalty is set to , meaning the maximum energy possible for a valid configuration for this problem instance, which results in the following function H(x):
The energy of the function is equal to , which is above the maximum energy possible for valid solutions of . This ensures that no invalid solution will have lower energy than any of the valid solutions. The technique demonstrated in this simplified example can be applied similarly to problems with more jobs and operations. The resulting QUBO will have more variables.
The same method is used to calculate the weights for other constraints as well. Generally, the Lagrange parameters are set individually for each constraint, depending on the magnitude of the objective function to be minimized and its relative strength to the constraints. In this case, to make sure that the cost function H(x) has the lowest energy for valid solutions, the weights of penalty functions f(x), g(x) and f(x) are all set equal to t, which is the maximum energy value of all valid solutions.
@@ 5 Experiments
In this Section, we implement the model and conduct experiments on a D-Wave quantum annealer (specifically, the Advantage System 4.1). As is common practice, the results of the experiment on a quantum annealer are compared with those obtained from SA. Firstly, the problem size (see 5.1) and solver parameters (see 5.2) are defined. The QUBO model is constructed using the PyQUBO Python library, which provides a high-level interface for formulating combinatorial optimization problems in QUBO form and supports solving them using various solvers, including the D-Wave quantum annealer [38]. We then present the results of the experiments in Subchapter 5.3 and discuss them in Subchapter 5.4.
@@ 5.1 Problem size and dataset
To analyze the performance of the D-Wave quantum annealer for the DFJSP, 6 problem instances are defined in such a way that the size of the smallest problem instance is 50 and the size of the largest one is 250. The instances in between are defined in even intervals of 50 (e.g., 100, 150, etc999Problem size is rounded for some instances (e.g. instance size 99 is rounded to 100).). The size of the problem instances refers to the number of binary variables xi,o,m,t necessary to represent the problem in QUBO form. The maximum size of the problem instances is dictated by the physical properties of the D-Wave QPU. The D-Wave Advantage System 4.1 has 5627 qubits. However, not all qubits are interconnected. When embedding the model variables onto the physical hardware graph of the QPU, in some cases, to represent a single logical variable set of physical qubits called chains are necessary, with couplings between physical qubits capable of realizing the correct interaction between logical variables [36]. This leads to embedding patterns that need more physical qubits than there are variables in the original problem.
The purpose of the experiment is to test the capacity and performance of the D-Wave quantum annealer for the DFJSP up to the largest problem that can be embedded into the QPU. Tbl. 3 summarizes the instances solved by QA, the number of qubits needed to embed the model onto the physical QPU, and the length of the longest chain in the embedding.
@@
Smaller problem instances require a lower number of qubits for embedding. It is also true that the length of the chains increases with problem size, meaning that with an increasing number of variables, it becomes more challenging to find embedding schemas. Therefore, longer chains are required to connect the variables through couplers. In this model, the largest problem instance that could be embedded was of size 250, which required more than 4000 qubits for embedding. For 300 variables, no embedding could be found.
The problem instances are formulated for 2 jobs, j and j, each job containing a minimum of 2 operations. Each operation can be processed on one machine or a set of two machines, with different processing times on each machine. To include shipping distance in the model, the sets of machines available for each operation include machines from different factories. The consideration of the shipping distance is essential for the processing of Contribution 1. The number of operations in each job, the available machines, and the processing time on each machine determine the number of binary variables xi,o,m,t necessary to represent the problem in QUBO form.
@@ 5.2 Solver parameters
There are numerous solver parameters101010https://docs.dwavequantum.com/en/latest/industrial_optimization/index_properties_parameters.html that can be adjusted when submitting problems to solvers. In this experiment, the default settings are kept for most of them, including the default annealing time of 20s. The parameters that are adjusted in the experiment are the number of reads and the chain strength.
Number of reads. In theory, very long annealing times should keep the system in a low energy state, in practice, noise due to hardware imperfections will push the state to a higher energy [5]. Therefore, short annealing times are preferred in order to take many samples from the energy distribution of the problem. The appropriate number of samples depends on the problem itself and its energy distribution, and is usually between 10 and 10,000 samples [5]. In this experiment, the number of reads (’num_reads’) is set to 1000. Generally, increasing the number of reads can increase the probability of obtaining better solutions, but improvements diminish due to some problem-dependent factors. For this model, increasing the number of reads significantly above 1000 does not appear to be advantageous. Sampling the problem instance with 200 variables 3000 times produced a solution with an energy of 34 and 2 violated constraints, which is worse than the energy of the best solution of 24 and 1 violated constraint obtained with 1000 reads. This is consistent with the results obtained by Carugno et al. [5], who found that increasing the number of reads from 1000 to 10000 did not have any beneficial effects.
Chain strength. For a solution to be consistent, the chain of qubits has to act as a single qubit, meaning that all the qubits in the chain must have the same values after annealing [5]. Broken chains show different values of the qubits at the end of the annealing process. Chains must be set strong enough to ensure that the qubits of the chain show the same value compared to other qubits. If the chain’s strength is too small, then the qubits will show different values. If the chain strength parameter is too high, it will impede the qubits in the chain from changing their state, causing the chain to behave like a separate entity that does not interact with other terms of the function.
We follow the D-Wave recommendation111111https://docs.dwavequantum.com/en/latest/quantum_research/embedding_guidance.html to determine the chain strength for each problem instance and start with two utility functions to estimate the chain strength:
@@
The first function returns a chain strength that is scaled to the problem bias range. The second function calculates a chain strength that attempts to compensate for the torque that would break the chain. Both methods should be viewed as a starting point for estimating the optimal chain strength, rather than a rule, since this parameter is problem-dependent. In the next step, the chain strength parameter found in step 1 should be doubled or tripled to check if this change has an impact on the percentage of broken chains.
@@
To find the best chain strength for the model, the model is run on the QPU for each problem instance, and the resulting energy and the percentage of broken chains in the best solution are analyzed. The intention is to find the lowest energy possible, while also having a low rate of broken chains to ensure a consistent solution. Fig. 3 shows the results of the measurements for the problem instance of size 150.
@@
The utility function scaled returns a chain strength of 32 with an energy of 69; the function uniform_torque returns a value of 309 with an energy of 966. To understand the relationship between these parameters, the range between these two chain strengths is divided into 11 intervals, and the energy and the percentage of broken chains in the best solutions for each of these points are measured. Additionally, the smallest chain strength value is divided by 2 to check whether a smaller parameter could improve the results. The results in the graph show that the lowest energy achieved corresponds to the next point in the interval from the point obtained with the scaled function. However, the percentage of broken chains corresponding to this value is 41%, which is high compared to values in the graph for higher chain strength values. To determine whether to accept a chain strength parameter of 39, an additional analysis of the solution’s consistency is necessary. Fig. 4 shows that, despite a higher number of broken chains, the embedding with a chain strength of 39 delivers a consistent solution without broken constraints. As the chain strength parameter increases, the number of broken constraints also increases.
@@
This procedure is repeated in the experiment for all problem instances to calculate the appropriate parameter for the chain strength. For all instances, the parameter that delivered the best solution with the lowest energy is one interval above the parameter estimated by the scaled function. Since the intention is to test the extent to which the QPU can handle large problems, this method is applied to problem instances of sizes 250 and 300. For problems of size 50 to 150, the solver can find consistent solutions with no broken constraints. For problem instances of size 200 and 250, the solutions with the lowest energy found had 1 and 2 broken constraints, respectively. For the instance of size 300, no embedding could be found.
@@
SA parameters. To solve problems with SA, the D-Wave Sampler class SimulatedAnnealingSampler of the D-Wave Ocean SDK is used. The sampler utilizes a classical heuristic SA algorithm that runs locally on the CPU121212Hardware configuration: 2x Intel Xeon Gold 6326 CPU 16Core, 2.9GHz 24M 185W Socket 4189, 4 x Samsung 64GB DDR4 ECC RDIMM 3200Mhz total 256GB RAM, 4 x HDD WD Gold Datacenter 1TB SATA, 1x SSD Samsung PM883 480GB SATA, 2 x 1 / 10 GB Ethernet 10 BaseT RJ-45 onboard.. For SA, the Central Processing Unit (CPU) time is measured, which is the amount of time that the CPU has spent solving the problem. The CPU time is compared to QPU access time.